The value of K such that the straight line (2x + 3y + 4) + K(6x - y +12) = 0 is parallel to y-axis is

Grade:12th

1 Answers

Arun
6 years ago
Equation of y axis
X = 0
Hence slope of line = infinity = 1/0
Now slope of other kine also should be infinity if they both are parallel.
Hence
Slope of other line,
(-2 -6k)/(3-k) = 1/0
3 - k = 0
Hence k = 3